# A class of charged and charged-Taub-NUT metrics in the presence of a massless scalar field and some of their astrophysical aspects

### v1

@@ 3 A Class of Charged Taub-NUT Metrics in the Presence of a Scalar Field
The Ehlers transformations for the Ernst potentials $\varepsilon$ and $\Phi$ are as follows
$\varepsilon^{\prime}=\frac{\varepsilon}{1+i\,c\,\varepsilon},\qquad\Phi^{\prime}=\frac{\Phi}{1+i\,c\,\varepsilon},$ (72)
where $c$ is a constant and real parameter. Using Eqs. (14), (39), and (40), we obtain the functions $\varepsilon$, $\Phi$, and $\chi$ as follows
@@
$\Phi^{\prime}=\frac{q\,r}{r^{2}+c^{2}\,(r-2\,m)^{2}}-i\,\frac{c\,q\,(r-2\,m)}{r^{2}+c^{2}\,(r-2\,m)^{2}}.$
@@
$F^{\prime}=\text{Re}(\varepsilon^{\prime})+\lvert\Phi^{\prime}\rvert^{2}=\frac{r^{2}-2\,m\,r+q^{2}}{r^{2}+c^{2}\,(r-2\,m)^{2}},$ (75)
@@
Also, considering the function $\Phi^{\prime}$ in Eq. (72), the functions $A_{t}^{\prime}$ and $\tilde{A}_{\phi}^{\prime}$ are determined by the following equations
$A_{t}^{\prime}=\text{Re}(\Phi^{\prime})=\frac{q\,r}{r^{2}+c^{2}\,(r-2\,m)^{2}},$ (76)
$\tilde{A}_{\phi}^{\prime}=\text{Im}(\Phi^{\prime})=-\,\frac{c\,(r-2\,m)^{2}}{r^{2}+c^{2}\,(r-2\,m)^{2}}.$
@@
×$\times\Big{[}\frac{dr^{2}}{r^{2}-2\,m\,r+q^{2}}+d\theta^{2}\Big{]}+\big{[}r^{2}+c^{2}\,(r-2\,m)^{2}\big{]}\,\sin^{2}\theta\,d\phi^{2}.$
@@
Using $\Box\varphi^{\prime}=0$ and metric (85), we obtain the scalar field $\varphi^{\prime}$ as follows
$\varphi^{\prime}=\sqrt{-\frac{\nu}{2}}\,\ln\Big{(}\frac{r-m-\sqrt{m^{2}+n^{2}-q^{2}}}{r-m+\sqrt{m^{2}+n^{2}-q^{2}}}\Big{)}.$ (88)
@@
3.1 Singularities of a Class of Charged Taub-NUT Metrics in the Presence of a Scalar Field
@@
3.2 AXIS SINGULARITIES AT $\theta=0$ AND $\theta=\pi$
@@
When $\nu=0$, the scalar field vanishes and Eq. (91) reduces to the charged Taub-NUT metric. It’s important to note that while the spacetime exhibits a singular behavior at $r=m\pm\sqrt{m^{2}+n^{2}-q^{2}}$, it remains regular at $r=0$. On the symmetry axis, however, i.e., at $\theta=0$ and $\theta=\pi$, the metric becomes singular due to the alteration of its signature, deviating from the usual (,+,+,+)(-,+,+,+)( - , + , + , + ) structure.
@@
This adjustment makes the axis at $\theta=0$ regular, allowing us to interpret $\phi$ as a periodic coordinate ranging from 00 to $2\pi$. However, the same singular behavior remains at $\theta=\pi$, manifesting as a conical defect.
@@
3.3 CLOSED TIMELIKE CURVES IN THE CHARGED TAUB-NUT METRICS IN THE PRESENCE OF A SCALAR FIELD
@@ 4 Gravitational lensing
4.1 Gravitational Lensing for a Class of Charged Metrics in the Presence of a Scalar Field
@@
[…]
@@
$\delta=-\,\int_{0}^{\pi}\,\int_{r_{0}}^{\infty}\,\partial_{r}\,\big{[}\mathcal{F}^{1-\frac{1}{2}\,(\gamma+\mu)}\;\mathcal{K}^{-\,\frac{\nu}{2}}\,\big{(}1+2\,r\,\frac{\mathcal{R}^{\prime}}{\mathcal{R}}+r\,\big{(}\frac{1}{2}-\gamma\big{)}\,\frac{\mathcal{F}^{\prime}}{\mathcal{F}}\big{)}\big{]}\,dr\,d\phi.$ (102)
@@
$\delta=$ $-\int_{0}^{\pi}\,\int_{r_{0}}^{\infty}\,\partial_{r}\,\big{\{}1-\frac{2\,m\,\gamma}{r}+\frac{1}{2\,r^{2}}\,\big{[}(q_{e}^{2}+q_{m}^{2})\big{(}4\,\mu^{2}+8\,\mu\,(\nu-1)+\nu\,(4\,\nu-7)+3\big{)}$ (103)
@@
If we choose $\mu=\gamma^{2}-\gamma$ and $\nu=1-\gamma^{2}$, the $\delta$ for the $\gamma$-metric is obtained as follows
@@
The difference in the deflection angle of light in the background of the charged FJNW and charged $\gamma$-metrics is as follows
@@
4.2 Gravitational Lensing in Charged Taub-NUT Metrics with a Scalar Field
In this subsection, we aim to derive the light deflection angle for gravitational lensing in the metrics given by (85). To achieve this, we follow the same approach as in subsection 4.1 and calculate the deflection angle for a class of charged Taub-NUT metrics in the presence of a scalar field, as outlined below
@@ 4.2 Gravitational Lensing in Charged Taub-NUT Metrics with a Scalar Field
5 Quasi normal modes
@@
5.1 QNMs of the Reissner-Nordström (RN) metric in the presence of a scalar field
@@
We apply the same approximations in the following calculation of the QNMs. Defining the QNMs as $Q=\Omega+i,\tilde{\Gamma}$, we first focus on obtaining the real part of the QNMs. In the eikonal limit, using the light ring method, $\Omega$ can be written as follows
$\Omega=\pm\,j\,\Omega_{\pm},$ (125)
where $\Omega_{\pm}$ is the frequency corresponding to the null rays of periodic orbits, and $j$ is a real number. Perturbations of massless waves in an axisymmetric spacetime can be expressed as the following eigenvalue superpositions
$e^{i\,(\Omega\,t-\iota\,\phi)}\,S_{\Omega\,j\,\iota\,s}(r,\theta).$ (126)
where $S$ and $\Omega$ are the spin and frequency of the waves, respectively, and $\iota$ and $j$ are the angular momentum parameters that satisfy
@@
$\Omega\gg 1/M,\qquad\lvert\iota\rvert=j\gg 1.$ (128)
From the calculations of QNMs associated with different black holes, it is determined that the frequency of perturbations ($\Omega$) is given by
$\Omega=\iota\,\frac{d\phi}{dt}=\pm\,j\,\Omega_{\pm}.$ (129)
To calculate $\Omega_{\pm}$, we first need to determine the radius of the null geodesic circles. For this, we use the following relations
@@
$\Omega_{\pm}=\frac{d\phi}{dt}=\frac{1}{3\,\sqrt{3}\,m}\,\big{(}1\pm\frac{q^{2}}{6\,m^{2}}\big{)}.$ (137)
Now, we will calculate $\tilde{\Gamma}$, which represents the divergence of null rays or the decay rate of the amplitude. To do this, we introduce a small perturbation to the coordinates $x^{\mu}=(t,r,\theta,\phi)=(t,r_{0},\frac{\pi}{2},\Omega_{\pm}t)$, thereby perturbing the null circular orbit in the equatorial plane
$r=r_{0}\,[1+\epsilon_{0}\,f_{r}(t)],\quad\phi=\Omega_{\pm}\,[t+\epsilon_{0}\,f_{\phi}(t)],\quad\ell=t+\epsilon_{0}\,f_{\ell}(t).$ (138)
@@
[…]
@@
$-\frac{\Omega_{\pm}}{\sqrt{-g}}\,\frac{\partial}{\partial\phi}\,\big{[}\sqrt{-g}\,(1-\epsilon_{0}\,f_{\ell}^{\prime})\big{]}.$
@@
$\frac{1}{\rho_{n}}\,\frac{d\rho_{n}}{d\ell}=-\epsilon_{0}\,\big{[}\frac{\partial(2\,f_{r}-f_{\ell}^{\prime})}{\partial t}+\frac{\partial(r_{0}\,f_{r}^{\prime})}{\partial r}+\Omega_{\pm}\,\frac{\partial(2\,f_{r}-f_{\ell}^{\prime})}{\partial\phi}\big{]}.$ (145)
@@
$\frac{d^{2}r}{d\ell^{2}}+\Gamma_{tt}^{r}\,\big{(}\frac{dt}{d\ell}\big{)}^{2}+\Gamma_{\phi\phi}^{r}\,\big{(}\frac{d\phi}{d\ell}\big{)}^{2}+\Gamma_{\theta\theta}^{r}\,\big{(}\frac{d\theta}{d\ell}\big{)}^{2}=0.$ (147)
@@
5.2 QNMs of the charged FJNW metric ($\nu=0$)
@@
$ds^{2}$ $=-\big{[}1-\frac{2\,m}{r}+\frac{q_{e}^{2}+q_{m}^{2}}{r^{2}}+\xi\,(1-\frac{2\,m}{r})\,\ln\big{(}1-\frac{2\,m}{r}\big{)}\big{]}\,dt^{2}+\frac{1}{1-2\,m/r}$ (154)
×$\times\big{[}1-\frac{q_{e}^{2}+q_{m}^{2}}{r^{2}\,(1-2\,m/r)}-\xi\,\ln\big{(}1-\frac{2\,m}{r}\big{)}\big{]}\,dr^{2}+r^{2}\,\big{[}1-\xi\,\ln\big{(}1-\frac{2\,m}{r}\big{)}\big{]}\,\big{(}d\theta^{2}+d\phi^{2}\big{)}.$
We can perform the same calculation as in subsection 5.1 to obtain the QNMs of the metric in Eq. (154) as follows
$Q$ $=\Omega+i\,\tilde{\Gamma}$ (155)
@@
5.3 QNMs of a class of charged Taub-NUT metrics in the presence of a scalar field
Here, we aim to obtain the QNMs of the metrics in Eq. (85) using the method introduced in subsection 5.1. To compute the QNMs, we first expand metric (85) to first order in $\nu$ and second order in $q$ and $n$, while ignoring terms of order $\nu,q$, $\nu,n$, $q^{2},n$, and $q,n^{2}$
@@ 5.3 QNMs of a class of charged Taub-NUT metrics in the presence of a scalar field
×$\times\big{[}1+\nu\,\ln\big{(}1+\frac{m^{2}}{r^{2}\,(1-2\,m/r)}\big{)}\big{]}\Big{)}\,dr^{2}+\Big{(}n^{2}+r^{2}\,\big{[}1+\nu\,\ln\big{(}1+\frac{m^{2}}{r^{2}\,(1-2\,m/r)}\big{)}\big{]}\Big{)}\,d\theta^{2}$

### v2

@@ III A Class of Charged Taub-NUT Metrics in the Presence of a Scalar Field
The Ehlers transformations for the Ernst potentials $\varepsilon$ and are as follows
[…]
where $c$ is a constant and real parameter. Using Eqs. (14), (39), and (40), we obtain the functions $\varepsilon$, , and $\chi$ as follows
@@
${}^{\prime}=\frac{q\,r}{r^{2}+c^{2}\,(r-2\,m)^{2}}-i\,\frac{c\,q\,(r-2\,m)}{r^{2}+c^{2}\,(r-2\,m)^{2}}.$
@@
$F^{\prime}=\text{Re}(\varepsilon^{\prime})+\lvert{}^{\prime}\rvert^{2}=\frac{r^{2}-2\,m\,r+q^{2}}{r^{2}+c^{2}\,(r-2\,m)^{2}},$ (75)
@@
Also, considering the function in Eq. (72), the functions $A_{t}^{\prime}$ and $\tilde{A}_{\phi}^{\prime}$ are determined by the following equations
$A_{t}^{\prime}=\text{Re}({}^{\prime})=\frac{q\,r}{r^{2}+c^{2}\,(r-2\,m)^{2}},$ (76)
$\tilde{A}_{\phi}^{\prime}=\text{Im}({}^{\prime})=-\,\frac{c\,(r-2\,m)^{2}}{r^{2}+c^{2}\,(r-2\,m)^{2}}.$
@@
$\times\Big[\frac{dr^{2}}{r^{2}-2\,m\,r+q^{2}}+d\theta^{2}\Big]+\big[r^{2}+c^{2}\,(r-2\,m)^{2}\big]\,\sin^{2}\theta\,d\phi^{2}.$
@@
Using $\square\varphi^{\prime}=0$ and metric (85), we obtain the scalar field $\varphi^{\prime}$ as follows
[…]
@@
III.1 Singularities of a Class of Charged Taub-NUT Metrics in the Presence of a Scalar Field
@@
III.2 AXIS SINGULARITIES AT $\theta=0$ AND $\theta=\pi$
@@
When $\nu=0$, the scalar field vanishes and Eq. (91) reduces to the charged Taub-NUT metric. It’s important to note that while the spacetime exhibits a singular behavior at $r=m\pm\sqrt{m^{2}+n^{2}-q^{2}}$, it remains regular at $r=0$. On the symmetry axis, however, i.e., at $\theta=0$ and $\theta=\pi$, the metric becomes singular due to the alteration of its signature, deviating from the usual (,+,+,+)(-,+,+,+) structure.
@@
This adjustment makes the axis at $\theta=0$ regular, allowing us to interpret $\phi$ as a periodic coordinate ranging from 0 to $2\pi$. However, the same singular behavior remains at $\theta=\pi$, manifesting as a conical defect.
@@
III.3 CLOSED TIMELIKE CURVES IN THE CHARGED TAUB-NUT METRICS IN THE PRESENCE OF A SCALAR FIELD
@@ IV Gravitational lensing
IV.1 Gravitational Lensing for a Class of Charged Metrics in the Presence of a Scalar Field
@@
$\delta=-\,\intop\nolimits_{0}^{\pi}\,\intop\nolimits_{r_{0}}^{\infty}\,\mathfrak{K}\,dS,$ (100)
@@
$\delta=-\,\intop\nolimits_{0}^{\pi}\,\intop\nolimits_{r_{0}}^{\infty}\,\partial_{r}\,\big[\mathcal{F}^{1-\frac{1}{2}\,(\gamma+\mu)}\;\mathcal{K}^{-\,\frac{\nu}{2}}\,\big(1+2\,r\,\frac{\mathcal{R}^{\prime}}{\mathcal{R}}+r\,\big(\frac{1}{2}-\gamma\big)\,\frac{\mathcal{F}^{\prime}}{\mathcal{F}}\big)\big]\,dr\,d\phi.$ (102)
@@
$\delta=$ $-\intop\nolimits_{0}^{\pi}\,\intop\nolimits_{r_{0}}^{\infty}\,\partial_{r}\,\big\{1-\frac{2\,m\,\gamma}{r}+\frac{1}{2\,r^{2}}\,\big[(q_{e}^{2}+q_{m}^{2})\big(4\,\mu^{2}+8\,\mu\,(\nu-1)+\nu\,(4\,\nu-7)+3\big)$ (103)
@@
If we choose $\mu=\gamma^{2}-\gamma$ and $\nu=1-\gamma^{2}$, the $\delta$ for the ZV metric is obtained as follows
@@
The difference in the deflection angle of light in the background of the charged FJNW and charged ZV metrics is as follows
@@
IV.2 Gravitational Lensing in Charged Taub-NUT Metrics with a Scalar Field
In this subsection, we aim to derive the light deflection angle for gravitational lensing in the metrics given by (85). To achieve this, we follow the same approach as in subsection IV.1 and calculate the deflection angle for a class of charged Taub-NUT metrics in the presence of a scalar field, as outlined below
@@ IV.2 Gravitational Lensing in Charged Taub-NUT Metrics with a Scalar Field
V Quasi normal modes
@@
V.1 QNMs of the Reissner-Nordström (RN) metric in the presence of a scalar field
@@
[…]
where ± is the frequency corresponding to the null rays of periodic orbits, and $j$ is a real number. Perturbations of massless waves in an axisymmetric spacetime can be expressed as the following eigenvalue superpositions
[…]
where $S$ and are the spin and frequency of the waves, respectively, and $\iota$ and $j$ are the angular momentum parameters that satisfy
@@
[…]
From the calculations of QNMs associated with different black holes, it is determined that the frequency of perturbations () is given by
[…]
To calculate ±, we first need to determine the radius of the null geodesic circles. For this, we use the following relations
@@
$_{\pm}=\frac{d\phi}{dt}=\frac{1}{3\,\sqrt{3}\,m}\,\big(1\pm\frac{q^{2}}{6\,m^{2}}\big).$ (137)
[…]
$r=r_{0}\,[1+\epsilon_{0}\,f_{r}(t)],\quad\phi={}_{\pm}\,[t+\epsilon_{0}\,f_{\phi}(t)],\quad\ell=t+\epsilon_{0}\,f_{\ell}(t).$ (138)
@@
$K^{\mu}=\frac{dx^{\mu}}{d\ell}=\big(1-\epsilon_{0}\,\frac{\partial\,f_{\ell}}{\partial\,t},\,\epsilon_{0}\,r_{0}\,\frac{\partial\,f_{r}}{\partial\,t},\,0,\,{}_{\pm}\,(1-\epsilon_{0}\,\frac{\partial\,f_{\ell}}{\partial\,t})\big).$ (139)
@@
$-\frac{{}_{\pm}}{\sqrt{-g}}\,\frac{\partial}{\partial\phi}\,\big[\sqrt{-g}\,(1-\epsilon_{0}\,f_{\ell}^{\prime})\big].$
@@
$\frac{1}{\rho_{n}}\,\frac{d\rho_{n}}{d\ell}=-\epsilon_{0}\,\big[\frac{\partial(2\,f_{r}-f_{\ell}^{\prime})}{\partial t}+\frac{\partial(r_{0}\,f_{r}^{\prime})}{\partial r}+{}_{\pm}\,\frac{\partial(2\,f_{r}-f_{\ell}^{\prime})}{\partial\phi}\big].$ (145)
@@
$\frac{d^{2}r}{d\ell^{2}}+{}_{tt}^{r}\,\big(\frac{dt}{d\ell}\big)^{2}+{}_{\phi\phi}^{r}\,\big(\frac{d\phi}{d\ell}\big)^{2}+{}_{\theta\theta}^{r}\,\big(\frac{d\theta}{d\ell}\big)^{2}=0.$ (147)
@@
V.2 QNMs of the charged FJNW metric ($\nu=0$)
@@
$ds^{2}$ $=-\big[1-\frac{2\,m}{r}+\frac{q_{e}^{2}+q_{m}^{2}}{r^{2}}+\xi\,(1-\frac{2\,m}{r})\,\ln\big(1-\frac{2\,m}{r}\big.)\big]\,dt^{2}+\frac{1}{1-2\,m/r}$ (154)
$\times\big[1-\frac{q_{e}^{2}+q_{m}^{2}}{r^{2}\,(1-2\,m/r)}-\xi\,\ln\big(1-\frac{2\,m}{r}\big.)\big]\,dr^{2}+r^{2}\,\big[1-\xi\,\ln\big(1-\frac{2\,m}{r}\big.)\big]\,\big(d\theta^{2}+d\phi^{2}\big).$
We can perform the same calculation as in subsection V.1 to obtain the QNMs of the metric in Eq. (154) as follows
[…]
@@
V.3 QNMs of a class of charged Taub-NUT metrics in the presence of a scalar field
Here, we aim to obtain the QNMs of the metrics in Eq. (85) using the method introduced in subsection V.1. To compute the QNMs, we first expand metric (85) to first order in $\nu$ and second order in $q$ and $n$, while ignoring terms of order $\nu,q$, $\nu,n$, $q^{2},n$, and $q,n^{2}$
@@ V.3 QNMs of a class of charged Taub-NUT metrics in the presence of a scalar field
$\times\big[1+\nu\,\ln\big(1+\frac{m^{2}}{r^{2}\,(1-2\,m/r)}\big.)\big]\Big)\,dr^{2}+\Big(n^{2}+r^{2}\,\big[1+\nu\,\ln\big(1+\frac{m^{2}}{r^{2}\,(1-2\,m/r)}\big.)\big]\Big)\,d\theta^{2}$